# Second Law of Thermodynamics without Einstein Relation

### v2

@@ Abstract
We now revisit additional classical-thermodynamic results demonstrating the relevance of the generalized temperature $\mathcal{T}$. More details regarding these consequences are found in the SM [67]. (i) Consider a process that starts and ends at the same protocol value, $\Lambda_{t_{\mathrm{f}}}=\Lambda_{0}$, and then allowed to relax to steady state. Since $\oint_{0}^{\infty}{\text{d}}t\langle\dot{S}_{t}\rangle=0$, we obtain the Clausius inequality, $\oint_{0}^{\infty}{\text{d}}t\langle\dot{Q}_{t}\rangle/(k_{\mathrm{B}}\mathcal{T})=-\oint_{0}^{\infty}{\text{d}}t\langle\dot{\Sigma}_{t}\rangle\leq 0$ [67]. (ii) Construct a Carnot-like cyclic engine consisting of an “isothermal” process (with constant $\mathcal{T}_{\mathrm{H}}$ and $\langle Q_{\mathrm{in}}\rangle>0$), an “adiabatic” process (in which the system “cools” down to $\mathcal{T}_{\mathrm{C}}$), followed by another isothermal process (with constant $\mathcal{T}_{\mathrm{C}}$, and $\langle Q_{\mathrm{out}}\rangle<0$), and an adiabatic process (where the system reaches $\mathcal{T}_{\mathrm{H}}$). The efficiency of such an engine is found as $\langle-W_{\mathrm{cycle}}\rangle/\langle Q_{\mathrm{in}}\rangle=1-\mathcal{T}_{\mathrm{C}}/\mathcal{T}_{\mathrm{H}}+k_{\mathrm{B}}\mathcal{T}_{\mathrm{C}}\langle\Sigma_{\mathrm{cycle}}\rangle/\langle Q_{\mathrm{in}}\rangle$, where $\langle\Sigma_{\mathrm{cycle}}\rangle$ is the mean IEP during a cycle [67]. Thus, the mean IEP quantifies the cycle’s irreversibility compared to a generalized Carnot efficiency, $1-\mathcal{T}_{\mathrm{C}}/\mathcal{T}_{\mathrm{H}}$. (See also Ref. [68] in the context of nonequilibrium engines.) (iii) Define the nonequilibrium free energy, $A_{t}=H_{t}-k_{\mathrm{B}}\mathcal{T}S_{t}$. It is easy to show that it bounds the extractable work in “isothermal” (constant $\mathcal{T}$) transformations, $\langle-\dot{W}_{t}\rangle=-(\langle\dot{A}_{t}\rangle+k_{\mathrm{B}}\mathcal{T}\langle\dot{\Sigma}_{t}\rangle)\leq-\langle\dot{A}_{t}\rangle$ [67].
@@
An interesting consequence of Eq. (15) relates to a particular class of nonequilibrium systems having multiple “temperatures”, e.g., a mixture of different types of particles, each of which is subjected to a different “thermostat” [69, 70]. These models may be relevant, for instance, for mixtures of passive and active particles [71, 72, 73, 74, 75]. They yield a temperature-like scalar which is a weighted average of the thermostats’ temperatures. Considering a diagonal but nonscalar $\mathbf{D}^{-1}\cdot\boldsymbol{\mu}$, with different inverse temperatures as its eigenvalues, we may view Eq. (15) as a generalization of such a weighted average. These systems may attain $\mathbf{J}=\mathbf{0}$, which results in a ‘$/0$ 0’ ambiguity for $\mathcal{T}$. However, perturbing different global parameters amounts to taking a perturbation $\mathbf{J}$ along different axes, thus picking up different eigenvalues (inverse temperatures) of $\mathbf{D}^{-1}\cdot\boldsymbol{\mu}$. We show a simplistic example in the SM [67].
@@
Conclusion.—We have analyzed a generic nonequilibrium Langevin system [Eq. (2)] which violates the ER [Eq. (1)]. This describes ubiquitous scenarios in driven and active materials since the relation between dissipation (mobility) and fluctuations (diffusivity) is a particular consequence of the equilibrium Boltzmann distribution [14]. We have demonstrated the far-reaching consequences of ER violation, which may be summarized as the detachment of informatic quantities (IH, IEP) from their physical thermodynamic counterparts (heat, dissipation) and the resulting invalidation of fluctuation theorems which rely on these correspondences.
@@
To restore the information-thermodynamics correspondence we have imposed the second law of thermodynamics directly and identified a generalized temperature $\mathcal{T}$ [Eq. (15)], which gives all the thermodynamically required properties: nonnegative dissipation, zero dissipation at equilibrium, linear thermodynamic stability, and the temperature being a macroscopic (integrated) property. The generalized temperature is model-independent — with a calorimetric measurement of the heat and path-sampling for the IH, one may find the system’s $\mathcal{T}$ with no required knowledge of the underlying dynamics. In particular, the variety of tools available for the estimation of IEP from trajectories (e.g., sampling marginal distributions [76, 77, 58] or the thermodynamic uncertainty relations [78, 79]) will be useful, since in our construction the average IEP remains connected to the average thermodynamic dissipation. The generalized temperature does not restore the various FRs [37, 39, 36, 65], which in the absence of the ER no longer dictate the statistics of thermodynamic properties [44].
@@
The three conditions which have been used to find $\mathcal{T}$, Eq. (15), do not uniquely determine it. Within linear irreversible thermodynamics [62, 61, 63], they do guarantee a single generalized temperature, as they impose the exact leading $\mathcal{O}(J^{2})$ term, $\langle\dot{R}_{t}\rangle=\int{\text{d}}\mathbf{x}(\mathbf{J}\cdot\mathbf{D}^{-1}\cdot\mathbf{J})/p$. This leading term is the complete dissipation functional when the Einstein relation holds. If one additionally requires this dissipation functional beyond linear theory, $\mathcal{T}$ is then uniquely determined by Eq. (15). The fact that $\mathcal{T}$ is the ratio between $\langle\dot{Q}_{t}\rangle$ (thermodynamics) and $\langle\dot{\Omega}_{t}\rangle$ (information) is then a useful consequence of this choice of dissipation functional.
@@ Acknowledgements.
We greatly benefited from discussions with David Dean and Yael Roichman. T.M. acknowledges funding from the Israel Science Foundation (Grant No. 1356/22).
@@
  • For reversible (near-equilibrium) processes, the Clausius inequality is an equality. Here, even if the Einstein relation does not apply, we find that at flux-free steady state, $\mathbf{J}=\mathbf{0}$, the generalized Clausius inequality becomes an equality. Moreover, it indicates that the flux-free state is stable (that is, for a small departure from equilibrium $\delta\mathbf{J}$, the dissipation is quadratic in it and hence much smaller). Lastly, based on various coarse graining techniques of the informatic entropy production (e.g., the thermodynamic uncertainty relations [78, 79]), if a positive estimate of $\langle\Sigma_{t}\rangle$ is found,one obtains a tighter bound than the usual Clausius inequality.

### v3

@@ Abstract
We now revisit additional classical-thermodynamic results demonstrating the relevance of the generalized temperature $\mathcal{T}$. More details regarding these consequences are found in the Supplemental Material [70]. (i) Consider a process that starts and ends at the same protocol value, $\Lambda_{t_{\mathrm{f}}}=\Lambda_{0}$, and then allowed to relax to steady state. Since $\oint_{0}^{\infty}{\text{d}}t\langle\dot{S}_{t}\rangle=0$, we obtain the Clausius inequality, $\oint_{0}^{\infty}{\text{d}}t\langle\dot{Q}_{t}\rangle/(k_{\mathrm{B}}\mathcal{T})=-\oint_{0}^{\infty}{\text{d}}t\langle\dot{\Sigma}_{t}\rangle\leq 0$ [70]. (ii) Construct a Carnot-like cyclic engine consisting of an “isothermal” process (with constant $\mathcal{T}_{\mathrm{H}}$ and $\langle Q_{\mathrm{in}}\rangle>0$), an “adiabatic” process (in which the system “cools” down to $\mathcal{T}_{\mathrm{C}}$), followed by another isothermal process (with constant $\mathcal{T}_{\mathrm{C}}$, and $\langle Q_{\mathrm{out}}\rangle<0$), and an adiabatic process (where the system reaches $\mathcal{T}_{\mathrm{H}}$). The efficiency of such an engine is found as $\langle-W_{\mathrm{cycle}}\rangle/\langle Q_{\mathrm{in}}\rangle=1-\mathcal{T}_{\mathrm{C}}/\mathcal{T}_{\mathrm{H}}+k_{\mathrm{B}}\mathcal{T}_{\mathrm{C}}\langle\Sigma_{\mathrm{cycle}}\rangle/\langle Q_{\mathrm{in}}\rangle$, where $\langle\Sigma_{\mathrm{cycle}}\rangle$ is the mean IEP during a cycle [70]. Thus, the mean IEP quantifies the cycle’s irreversibility compared to a generalized Carnot efficiency, $1-\mathcal{T}_{\mathrm{C}}/\mathcal{T}_{\mathrm{H}}$. (See also Ref. [71] in the context of nonequilibrium engines.) (iii) Define the nonequilibrium free energy, $A_{t}=H_{t}-k_{\mathrm{B}}\mathcal{T}S_{t}$. It is easy to show that it bounds the extractable work in isothermal (constant $\mathcal{T}>0$) transformations, $\langle-\dot{W}_{t}\rangle=-(\langle\dot{A}_{t}\rangle+k_{\mathrm{B}}\mathcal{T}\langle\dot{\Sigma}_{t}\rangle)\leq-\langle\dot{A}_{t}\rangle$ [70].
@@
An interesting consequence of Eq. (15) relates to a particular class of nonequilibrium systems having multiple “temperatures”, e.g., a mixture of different types of particles, each of which is subjected to a different “thermostat” [72, 73]. These models may be relevant, for instance, for mixtures of passive and active particles [74, 75, 76, 77, 78]. They yield a temperaturelike scalar that is a weighted average of the thermostats’ temperatures. Considering a diagonal but nonscalar $\mathbf{D}^{-1}\cdot\boldsymbol{\mu}$, with different inverse temperatures as its eigenvalues, we may view Eq. (15) as a generalization of such a weighted average. These systems may attain $\mathbf{J}=\mathbf{0}$, which results in a “$/0$ 0” ambiguity for $\mathcal{T}$. However, perturbing different global parameters amounts to taking a perturbation $\mathbf{J}$ along different axes, thus picking up different eigenvalues (inverse temperatures) of $\mathbf{D}^{-1}\cdot\boldsymbol{\mu}$. We show a simplistic example in the Supplemental Material [70].
@@
Conclusion.—We have analyzed a generic nonequilibrium Langevin system [Eq. (2)] that violates the ER [Eq. (1)]. This describes ubiquitous scenarios in driven and active materials since the relation between dissipation (mobility) and fluctuations (diffusivity) is a particular consequence of the equilibrium Boltzmann distribution [14]. We have demonstrated the far-reaching consequences of ER violation, which may be summarized as the detachment of informatic quantities (IH, IEP) from their physical thermodynamic counterparts (heat, dissipation) and the resulting invalidation of fluctuation theorems which rely on these correspondences.
@@
To restore the information-thermodynamics correspondence we have imposed the second law of thermodynamics directly and identified a generalized temperature $\mathcal{T}$ [Eq. (15)], which gives all the thermodynamically required properties: non-negative dissipation, zero dissipation at equilibrium, linear thermodynamic stability, and the temperature being a macroscopic (integrated) property. The generalized temperature is model-independent—with a calorimetric measurement of the heat and path sampling for the IH, one may find the system’s $\mathcal{T}$ with no required knowledge of the underlying dynamics. In particular, the variety of tools available for the estimation of IEP from trajectories (e.g., sampling marginal distributions [79, 80, 60] or the thermodynamic uncertainty relations [81, 82]) will be useful, since in our construction the average IEP remains connected to the average thermodynamic dissipation. The generalized temperature does not restore the various FRs [38, 40, 37, 68], which in the absence of the ER no longer dictate the statistics of thermodynamic properties [45].
@@
The three conditions that have been used to find $\mathcal{T}$, Eq. (15), do not uniquely determine it. Within linear irreversible thermodynamics [64, 63, 65], they do guarantee a single generalized temperature, as they impose the exact leading $\mathcal{O}(J^{2})$ term, $\langle\dot{R}_{t}\rangle=\int{\text{d}}\mathbf{x}(\mathbf{J}\cdot\mathbf{D}^{-1}\cdot\mathbf{J})/p$. This leading term is the complete dissipation functional when the Einstein relation holds. If one additionally requires this dissipation functional beyond linear theory, $\mathcal{T}$ is then uniquely determined by Eq. (15). The fact that $\mathcal{T}$ is the ratio between $\langle\dot{Q}_{t}\rangle$ (thermodynamics) and $\langle\dot{\Omega}_{t}\rangle$ (information) is then a useful consequence of this choice of dissipation functional.
@@ Acknowledgements.
We greatly benefited from discussions with David Dean and Yael Roichman. T.M. acknowledges funding from the Israel Science Foundation (Grant No. 1356/22). G.A. and H.D. acknowledge funding from the Israel Science Foundation (Grant No. 1611/24).
@@
  • For reversible (near-equilibrium) processes, the Clausius inequality is an equality. Here, even if the Einstein relation does not apply, we find that at flux-free steady state, $\mathbf{J}=\mathbf{0}$, the generalized Clausius ineqsuality becomes an equality. Moreover, it indicates that the flux-free state is stable (that is, for a small departure from equilibrium $\delta\mathbf{J}$, the dissipation is quadratic in it and hence much smaller). Lastly, based on various coarse graining techniques of the informatic entropy production (e.g., the thermodynamic uncertainty relations [81, 82]), if a positive estimate of $\langle\Sigma_{t}\rangle$ is found,one obtains a tighter bound than the usual Clausius inequality.